# Classifying Infinity Topoi via weighted limits

### v1

@@ 6. The Object Classifier
$((TOP^{alg})^{/\mathcal{B}})^{th}\simeq Top_{/\mathcal{B}}^{alg}$
@@
By the first paragraphs of [MW25, Section 2.3.4], $Top_{\mathcal{B}}$ has a cotensoring over $Cat_{\infty}$. Thus, we have a chain of equivalences for two $\mathcal{B}$-topoi $A,B$ and a finite simplicial set $K$
@@
for each finite simplicial set $K$, where the vertical maps exhibit the universal property of the tensor product. The top horizontal map is thus an equivalence. We conclude using [Lur09a, Proposition 2.2.5.7], that
@@
is an equivalence of $\infty$-categories. The fact that $\Gamma$ is an equivalecne follows using the description of equivalences of $\infty$-bicategories from [GHL21, Remark 1.2.14].
@@
where each $g_{!},h_{!}$ are the right adjoints of $g^{!},h^{!}$.
@@
where $\Gamma$ is the global sections functor and the horizontal composites are forgetful functors. The left vertical map is an equivalence by Theorem 3.12. Thus, the bottom horizontal map is corepresentable.
@@ 7. Classifying topoi
The definition above is intentionally too general and has a very simple paradigm in mind: we identify a theory with its semantic trace. To explain our intution, and present the first interesting example of theory and classifying topos, let us restrict to the special case in which $\mathcal{B}$ is the topos of spaces $\mathcal{S}$. Through the whole section, we will make this restriction several times as this is simultaneously the most interesting case, and the case in which it is easier and most paradigmatic to present the examples.
@@
Example 7.10 ((Geometric) Sketches).
Lawvere theories are an extremely restricted framework (and yet so powerful). In [CGMC25], the authors present a generalization of the classical theory of sketches to the context of $\infty$-category. We say that a sketch $\Sigma$ in the sense of [CGMC25, Def 3.1] is geometric if its limit part consists of finite diagrams. Now, let $\Sigma$ be a geometric sketch, then we have an associated prestack
@@
mapping a topos $\mathcal{E}$ to the accessible category of models of $\Sigma$ in $\mathcal{E}$ in the sense of [CGMC25, page 9]. The construction is funtorial because the sketch is geometric, indeed inverse images will preserve the colimit part because they are cocontinuous and the finite (!) limit diagram because they are lex. In the aforementioned paper the authors show that a very vast number of constructions are sketchable, and it is easy to see that most of them (if not all) are given by geometric sketches. Notice that geometric sketches generalise Lawvere theories on the spot, and in the classic case of $1$-dimensional topos theory this technology is equivalent to geometric logic.
@@
    We the sketch $\Sigma$ has no limit cones, nor colimit cocones, then it is easy to see that we have a classifier. This follows directly from Diaconescu, taking the free category with finite limits over $\mathcal{A}$.
@@
    Now assume for simplicity that $\Sigma$ has precisely one cone. In practise such structure can be understood as a couple of diagrams $D,\mathsf{const}_{x}:I\rightrightarrows\mathcal{A}$, where one of the diagrams is constant on a certain object $c$, together with a natural transformation $\rho:D\Rightarrow\mathsf{const}_{x}$ between them. Now, we can identify both $\mathcal{A}$ and $I$ with their associated sketch of $A$ and $I$-diagrams, obtaining two sketches $\Sigma^{\mathcal{A}}$ and $\Sigma^{I}$. Then, we obtain the following diagram,
@@
To finish off our discussion we recall definitions taken by [Joh02, B4.2] which invest on this precise intuition.
@@
  • Let $(\mathbb{T},U_{\mathbb{T}}),(\mathbb{S},U_{\mathbb{S}})$ geometric constructs (over the same sorting), a morphism of geometric construct is a natural transfromation like below,

### v2

@@ 6. The Object Classifier
$((TOP^{alg})^{/\mathcal{B}})^{th}\simeq Top_{/\mathcal{B}}^{alg}.$
@@
By the first paragraphs of [MW25, Section 2.3.4], $Top_{\mathcal{B}}$ has a cotensoring over $Cat_{\infty}$. Thus, using Lemma 1.18, we have a chain of equivalences for two $\mathcal{B}$-topoi $A,B$ and a finite $\infty$-category $K$
@@
for each finite $\infty$-category $K$, where the vertical maps exhibit the universal property of the cotensor product. The top horizontal map is thus an equivalence. We conclude using [Lur09a, Proposition 2.2.5.7], that
@@
is an equivalence of $\infty$-categories. The fact that $\Gamma$ is an equivalence follows using the description of equivalences of $\infty$-bicategories from [GHL21, Remark 1.2.14].
@@
where each $g^{!},h^{!}$ are the right adjoints of $g_{!},h_{!}$.
@@
where $\Gamma$ is the global sections functor and the horizontal composites are forgetful functors. The left vertical map is an equivalence by Lemma 6.4. Thus, the bottom horizontal map is corepresentable.
@@ 7. Classifying topoi
The definition above is intentionally too general and has a very simple paradigm in mind: we identify a theory with its semantic trace.
This notion of theory and classifying topos is free-styled in [ABFJ22, Section 5], where the authors present some of our examples below. Our discussion will upgrade their understanding by relating it to the notion of geometric sketch in the next subsection. Our treatment is indeed much less hand-wavy in assessing its expressive power and identifies clearly the range of expression of classifying topoi.
To explain our intuition, and present the first interesting example of theory and classifying topos, let us restrict to the special case in which $\mathcal{B}$ is the topos of spaces $\mathcal{S}$. Through the whole section, we will make this restriction several times as this is simultaneously the most interesting case, and the case in which it is easier and most paradigmatic to present the examples.
@@
7.1. Geometric sketches
We shall start with a short digression on the theory of sketches. The community of category theory and categorical logic introduced sketches as compact presentations of theories. Their intention is to generalize the notion of Lawvere theory to a much broader context, taking the blueprint of functorial semantics to its limit. There are many classical introductions to the notion of sketch in $1$-dimensional category theory, and the interested reader who wants to dive into the topic from a perspective closer to that of this paper can check [AR94, 2.F] and [Bor94, Vol II, 5.6].
Sketches are an extremely broad framework that encompasses almost every possible (predicate-style) logical system [AR94, Chap. 5]. As an evidence of the latter statement, and more tightly related to the topic of this paper, classically it is shown that sketches can be used as an alternative – purely categorical – framework to encode geometric logic, via the notion of geometric sketch (see [Bor94, Vol III, 4.3].
In [CGMC25], the authors present a generalization of the classical theory of sketches to the context of $\infty$-categories. We refer to the paper for a crash course on the topic, especially for a collection of examples. Notice that [CGMC25] proves that every example that we have presented in the previous part of the section is indeed given by the category of models of a sketch, and thus this subsection subsumes the previous discussion.
Definition 7.10.
We say that a sketch $\Sigma$ in the sense of [CGMC25, Def 3.1] is geometric if its limit part consists of finite diagrams. No restrictions on its colimit part.
Example 7.11 (The prestack associated to a geometric sketch).
Now, let $\Sigma$ be a geometric sketch, then we have an associated prestack
@@ 7.1. Geometric sketches
mapping a topos $\mathcal{E}$ to the accessible category of models of $\Sigma$ in $\mathcal{E}$ in the sense of [CGMC25, page 9]. The construction is functorial because the sketch is geometric, indeed inverse images will preserve the colimit part because they are cocontinuous and the finite (!) limit diagram because they are lex.
Remark 7.12.
In the aforementioned paper the authors show that a vast number of constructions are sketchable, and it is easy to see that most of them (if not all) are given by geometric sketches. Notice that geometric sketches generalize Lawvere theories on the spot, and in the classic case of $1$-dimensional topos theory this technology is equivalent to geometric logic.
Theorem 7.13.
If a theory has a classifying topos, then it is the prestack of models of a (large) geometric sketch.
Proof.
Let $\mathbb{T}$ be the theory and $\mathcal{B}[\mathbb{T}]$ be its classifying topos. We shall study only the case $\mathcal{S}[\mathbb{T}]$ for simplicity. Now, we equip $\mathcal{S}[\mathbb{T}]$ with a sketch structure, which we shall denote by $\Sigma[\mathbb{T}]$. In the colimit part we put all colimit cocone diagrams, and in the limit part we put all finite limit cone diagrams. It goes without saying that this is a geometric sketch. Now, it is clear that $\textbf{Mod}(\Sigma[\mathbb{T}],-)\simeq\textbf{Cocontlex}(\mathcal{S}[\mathbb{T}],-)$, where the latter is the category of cocontinuous and finite limit preserving functors. Observe that by the Adjoint Functor Theorem we have the following chain of equivalences, which finishes the proof.
$\textbf{Mod}(\Sigma[\mathbb{T}],-)\simeq\textbf{Cocontlex}(\mathcal{S}[\mathbb{T}],-)\stackrel{{\scriptstyle\text{AFT}}}{{\simeq}}TOP_{\mathcal{S}}(-,\mathcal{S}[\mathbb{T}])\simeq\mathbb{T}(-).$
Theorem 7.14.
If a theory has a classifying topos, then it is the prestack of models of a small geometric sketch.
Proof.
Let $\mathbb{T}$ be the theory, $\mathcal{B}[\mathbb{\mathbb{T}}]$ be its classifying topos and $\Sigma[\mathbb{T}]$ the large geometric sketch from the previous theorem. We shall provide a small geometric sketch $\Lambda[\mathbb{T}]$ and a morphism of sketches $i:\Lambda[\mathbb{T}]\to\Sigma[\mathbb{T}]$ that is a Morita equivalence, in the sense that the induced morphism below induces a natural equivalence of prestacks. This will finish the proof.
$\textbf{Mod}(\Sigma[\mathbb{T}],-)\simeq\textbf{Mod}(\Lambda[\mathbb{T}],-).$
To do so, consider any site $(C,\Xi)$ of definition (whose underlying category is lex) for $\mathcal{B}[\mathbb{T}]$ in the sense of [ABFJ22, 4.3.8]. It follows from [ABFJ22, 4.3.2 and the discussion right below] that we can assume $\Xi$ to consist of maps $\delta:P\to yc$ where $P$ is a presheaf, and $yc$ is a representable. Now, we can turn $\delta$ into a cocone between the diagram induced by the category of elements $\pi_{P}:\text{Elts}(P)\to Psh(C)$ and the constant diagram $const_{c}:\text{Elts}(P)\to Psh(C)$. These will be our colimit cocones in $\Lambda[\mathbb{T}]$, while its limit cones will be all finite limiting cones. It is easy to see that the inclusion $i:\Lambda[\mathbb{T}]\to\Sigma[\mathbb{T}]$ is a morphism of sketches with these definitions, and that it has the required properties, again by (the proof of) [ABFJ22, 4.3.10]. ∎
Remark 7.15.
The two proofs above are largely influenced by [DLL24, Sections 4.1 and 4.2]. Here, we adapted the arguments in that paper to the technology and language of [ABFJ22].
7.1.1. From geometric sketches to classifying topoi
At this point we have learned that if a theory has a classifying topos then it was actually the prestack of models of a geometric sketch. In the next part of the subsection we will prove the converse implication, i.e. every geometric sketch admits a classifying topos.
@@ 7.1.1. From geometric sketches to classifying topoi
    When the sketch $\Sigma$ has no limit cones, nor colimit cocones, then it is easy to see that we have a classifier. This follows directly from Diaconescu, taking the free category with finite limits over $\mathcal{A}$.
@@
    Now assume for simplicity that $\Sigma$ has precisely one cone. In practice such structure can be understood as a couple of diagrams $D,\mathsf{const}_{x}:I\rightrightarrows\mathcal{A}$, where one of the diagrams is constant on a certain object $c$, together with a natural transformation $\rho:D\Rightarrow\mathsf{const}_{x}$ between them. Now, we can identify both $\mathcal{A}$ and $I$ with their associated sketch of $A$ and $I$-diagrams, obtaining two sketches $\Sigma^{\mathcal{A}}$ and $\Sigma^{I}$. Then, we obtain the following diagram,
@@
To finish off our discussion we recall definitions taken by [Joh02, B4.2] which rest on this precise intuition.
@@
  • Let $(\mathbb{T},U_{\mathbb{T}}),(\mathbb{S},U_{\mathbb{S}})$ geometric constructs (over the same sorting), a morphism of geometric constructs is a natural transformation as below: